Considering the traits round (dominant) versus wrinkled (recessive) and yellow (dominant) versus green (recessive), determine the genotypes of the parental plants by analyzing the phenotypes of their offspring.

Wrinkled yellow X round yellow = 6/16 wrinkled yellow, 2/16 wrinkled green, 6/16 round yellow, 2/16 round green.
Answer #1

Answer:

Wrinkled yellow = rrYy

round yellow = RrYy

Explanation:

Wrinkled yellow (rrYy) X (RrYy) round yellow ----Parents

RY

Ry

rY

ry

rY

RrYY (rournd, yellow)

RrYy (rournd, yellow)

rrYY(wrinkled, yellow)

rrYy (wrinkled, yellow)

ry

RrYy (rournd, yellow)

Rryy (rournd, green)

rrYy (wrinkled, yellow)

rryy (wrinkled, green)

Round, yellow = 3/8

Round, green = 1/8

Wrinkled, yellow = 3/8

Wrinkled, green = 1/8

The obtained proportions are equal to given ratios as below:

6/16 wrinkled yellow = 6/16 = 3/8

2/16 wrinkled green = 1/8

6/16 round yellow = 3/8

2/16 round green = 1/8